Solve quadratic equation using the following steps x2 =12x -64 = 0
a) Move the constant term to the right side of the equation
b) Multiply each term in the equation by four time the coefficient of the x2 term
c) Square the coefficient of the original x term and add it to both sides of the equation
d) Take the square root of both sides
e) Set the left side of the equation equal to the negative square root on the right side of the equation and solve for x

Answer by mananth(16946)

x^2+12x-64=0
x^2+12x=64
add 36
x^2+12x+36=64+36
(x+6)^2=100
(x+6)^2=10^2
take the square root
x+6=+/-10
x=10-6
x=4
OR
x= -10-6
x=-16